Fermi National Laboratory


Accelerator Update

Monday June 28
The day shift began with Operations monitoring store 3597 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and the Meson experiment MCenter taking beam.

At 9:40 AM, MCenter called saying they were finished with beam.

A Main Injector (MI) expert began a slipstacking study at 10 AM.

Operations and experts stashed antiprotons into the Recycler at 10:37 AM. Pbar resumed stacking at 11:16 AM.

Booster experts began a collimator study at 1:34 PM. The study ended at 4:15 PM.

The MI slipstacking study ended at 5:15 PM.

At 5:54 PM, a Linac RF station (LRF1) tripped off. It reset okay.

At 7:07 PM, a Booster RF station (BRF12) tripped off. It reset okay.

At 7:39 PM, a Linac RF station (KRF5) tripped off. It reset okay.

Operations prepared the Meson MTest experiment at 8:41 PM to take beam.

At 8:57 PM, a Linac RF station (KRF5) tripped off. It reset okay.

Operations established beam to MTest at 9:03 PM.

At 9:29 PM, a Linac RF station (LRF2) tripped off. It reset okay.

At 9:44 PM, a Booster RF station (BRF4) tripped off. It reset okay.

Tuesday June 29
The midnight shift began with Operations monitoring store 3597 and the stash, with Pbar stacking, and with MiniBooNE and MTest taking beam.

Operations established beam to MCenter at 12:07 AM.

At 1:13 AM, a Booster RF station (BRF12) tripped off. It reset okay.

At 5:57 AM, a Booster RF station (BRF16) tripped off. It reset okay.

At 8 AM, the MCenter experimenters accidentally dropped their interlocks and then decided that they were finished with beam for a while.

Operations terminated store 3597 at 8:02 AM.

At 8:50 AM, a Booster RF station (BRF12) tripped off. It reset okay.

TeV experts began a study period at 9:09 AM.

At 9:27 AM, a Linac RF station (LRF1) tripped off. It reset okay.

Operations established beam to MTest at 9:34 AM.

At 10:14 AM, a Linac RF station (KRF5) tripped off for the first of three times in four minutes. It reset okay each time.

At 10:27 AM, the Neutron Therapy Facility (NTF) began running some study beam. Unfortunately, it was causing serious Linac beam problems. A Linac expert began tuning the beam line.

At 11:51 AM, a Booster RF station (BRF12) tripped off. It reset okay.

Operations put the TeV into shot setup at 1:33 PM.

Operations had loaded the TeV with protons and antiprotons (from Recycler and Pbar) and started ramping when, at 3:41 PM, the TeV suffered a sector E1 quench. TeV experts investigated and later discovered that the MI antiprotons had a large longitudinal emittance that turned into DC antiprotons in the TeV.

Pbar resumed stacking at 3:51 PM.

Operations prepared the D-Zero (D0) collision hall for access at 4:30 PM.

The Collider Detector at Fermilab (CDF) accessed their collision hall at 5:30 PM.

CDF and D0 came out of access at 6:30 PM.

At 9:43 PM, a TeV sextupole power supply (T:SF) tripped off on an overvoltage indication during a TeV dry squeeze study. Operators conducted a hipot due to the suspicious nature of this trip. The investigation continued, but the magnet might have ground faulted.

Operators prepared the TeV for access to search sector C3 for a ground fault at 11:31 PM.

Operations established beam to MCenter at 11:45 PM.

Wednesday June 30
The midnight shift began with operators and TeV experts investigating a TeV ground fault, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 12:10 AM, a Booster RF station (BRF11) tripped off. It would not reset. Operators bypassed the station.

A TeV expert reported at 12:49 AM that they found the ground fault at TeV sector A2. He also cleaned the sextupole power lead connections.

Everyone came out of the TeV and then Operations began the TeV turn-on at 1:37 AM.

At 2:43 AM, Operations put the TeV turn-on on hold due to a mechanical dump failure at sector B3. A TeV expert came in at 3:46 AM to investigate.

The Cryo Coordinator reported at 4:05 AM that the cryo system technicians were turning on other cryo compressors to compensate for a power outage at the B0/BA compressor room.

At 4:45 AM, the Duty Electrician reset the B0 breaker. When cryo system techs at CHL turned on the B0 cryo compressor it sparked and smoked, and caused a fire alarm.

The TeV expert made a new cable for the mechanical dump and the fault cleared at 7 AM. Operations restarted the TeV turn-on.

At 7:22 AM, a MI RF station (MIRF9) tripped off. It reset okay.

MCenter finished using beam at 7:27 AM.

At 7:43 AM, a TeV expert hipotted the sextupole magnet (T:SF) again and found it bad. He said another access was needed.

Operations prepared the TeV for an access at 8:13 AM.

Plans for Today
The plan for today is to further investigate the TeV SF magnet problem and bypass it if the magnet is ground faulted. During the repair time, MI experts will be allowed to conduct slipstacking studies, and NuMI experts will test a special magnet. Operations will put the TeV in shot setup sometime around 1 PM or as soon the TeV work gets finished. If all goes well, Recycler will stash antiprotons after the store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/30/2004   email Fermilab